# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v1

@@ A.1 Unlearning Baselines
We divide unlearning baselines into two categories based on whether they use the retain set $(\mathcal{D}_{r})$: (i) baselines without $\mathcal{D}_{r}$, and (ii) baselines with $\mathcal{D}_{r}$. The details of these algorithms are as follows.
Unlearning w/o $\mathcal{D}_{r}$. In this section, baselines aim to remove information using only $\mathcal{D}_{f}$ without referencing the retain set. While this reduces time complexity and privacy risk, it often sacrifices some accuracy on retain set.
@@
Pseudo Labeling (PL [7]). PL ensures that $\mathcal{D}_{f}$ is unlearned in a manner that minimizes noise. For each sample $x_{i}\in\mathcal{D}_{f}$, the original model $\theta_{o}$ assigns a pseudo-label $y_{i}^{\prime}$, representing the class with the highest predicted probability after excluding forget set. The unlearning objective is defined as:
@@ A.2 Representation-Based Metrics
where $K$ and $L$ represent the Gram matrices of the two feature representations, and HSIC denotes the Hilbert-Schmidt independence criterion. By applying CKA, we measure the similarity between the feature representations of $\theta_{u}$ (the unlearned model) and $\theta_{o}$ (the original model), $\theta_{r}$ (the retrained model) across various downstream datasets. Note that higher CKA values indicate a greater degree of alignment, suggesting that the unlearned model retains the desired feature representations after unlearning.
$k$-Nearest Neighbors ($k$-NN). To evaluate the transferability of the unlearned model, we leverage $k$-NN. Feature embeddings are extracted from the unlearned model, $\theta_{u}$, by feeding it the downstream data. These embeddings are then split, with 80% used to train a $k$-NN classifier and the remaining 20% used for testing. For each test sample, the classifier assigns a label based on the majority vote of its $k$ nearest neighbors in the training set. The $k$-NN accuracy is computed as the proportion of correctly classified test samples. This evaluation provides insights into whether $\theta_{u}$ preserves meaningful, transferable feature embeddings that are effective for downstream tasks.
@@ B.1 Training and Retraining Details
For the experiments, both the original and retrained models are required. In the case of ResNet-50, we trained the model from scratch. For ConvNeXt-T and Swin-T, we initialized them with ImageNet-21K pretrained weights [19]. Detailed configuration information is provided in Table 2.
@@ B.2 Unlearning Details
We refer to the hyperparameters from the original studies. However, because most prior works proposed and tested these methods primarily on smaller datasets (e.g., CIFAR-10, CIFAR-100), critical implementation details are not fully specified when removing multiple classes from large-scale datasets. To address this limitation, we perform a grid search for each method to identify the optimal learning rate and number of epochs. Specifically, we retrain the model (excluding the forget set $\mathcal{D}_{f}$) using the same learning rate and epochs as in the original training for the Retrain method. We limit FT at 40 epochs to mitigate excessive time costs. For PL, GA, RL, and SalUn, we find that strong unlearning can be achieved within 10 epochs on large-scale data. Finally, considering the balance of logit-based and representation-based metrics, we fix the maximum epochs for DUCK, CU, SCAR, and SCRUB at 15, 80, 20, and 90, respectively. These upper limits are determined first under the scenario of Random 100 class unlearning. For Random $k$ class experiments ($k\in\{100,200,300\}$), we choose the actual number of epochs while respecting these maximum values, guided by the H-LR score. For Top $k$ class unlearning, we reuse the hyperparameter configurations established in the corresponding Random $k$ setting.
@@ Appendix C Additional Experiments
Table 3 provides an extended comparison of the baselines using alternative backbones, specifically ConvNeXt-T [25] and Swin-T [24]. We only compare model-agnostic methods.

### v2

@@ A.1 Unlearning Baselines
We divide unlearning baselines into two categories based on whether they use retain set ($\mathcal{D}_{r}$) or not: (i) baselines without $\mathcal{D}_{r}$, and (ii) baselines with $\mathcal{D}_{r}$. The details of these algorithms are as follows.
Unlearning w/o $\mathcal{D}_{r}$. In this section, baselines aim to remove information using only $\mathcal{D}_{f}$ without referencing retain set. While this reduces time complexity and the privacy risk, it often sacrifices some accuracy on retain set.
@@
Pseudo Labeling (PL) [7]. PL ensures that $\mathcal{D}_{f}$ is unlearned in a manner that minimizes noise. For each sample $x_{i}\in\mathcal{D}_{f}$, the original model $\theta_{o}$ assigns a pseudo-label $y_{i}^{\prime}$, representing the class with the highest predicted probability after excluding forget set. The unlearning objective is defined as:
@@ A.2 Representation-Based Metrics
where $K$ and $L$ represent the Gram matrices of the two feature representations, and HSIC denotes the Hilbert-Schmidt independence criterion. By applying CKA, we measure the similarity between the feature representations of $\theta_{u}$ (the unlearned model) and $\theta_{o}$ (the original model), $\theta_{r}$ (the retrained model) across various downstream datasets. Note that higher CKA values indicate a greater degree of alignment, suggesting that the unlearned model retains the desired feature representations after unlearning. We use the average pooling layer of ResNet-50 to extract feature representations.
$k$-Nearest Neighbors ($k$-NN). To evaluate the transferability of the unlearned model, we leverage $k$-NN. Feature embeddings are extracted from the unlearned model, $\theta_{u}$, by feeding it the downstream data. These embeddings are then split, with 80% used to train a $k$-NN classifier and the remaining 20% used for testing. For each test sample, the classifier assigns a label based on the majority vote of its $k$ nearest neighbors in the training set. The $k$-NN accuracy is computed as the proportion of correctly classified test samples. This evaluation provides insights into whether $\theta_{u}$ preserves meaningful, transferable feature embeddings that are effective for downstream tasks. We set $k=5$ in the $k$-Nearest Neighbors ($k$-NN) classifier for all experiments.
@@ B.1 Training and Retraining Details
For the experiments, both the original and retrained models are required. In the case of ResNet-50, we trained it from scratch. For ConvNeXt-T and Swin-T, we initialized them with ImageNet-21K pretrained weights [19]. Detailed configuration information is provided in Table 2.
@@ B.2 Unlearning Details
We refer to the hyperparameters from the original studies. However, because most prior works proposed and tested these methods primarily on smaller datasets (e.g., CIFAR-10, CIFAR-100), critical implementation details are not fully specified when removing multiple classes from large-scale datasets. To address this limitation, we perform a grid search for each method to identify the optimal learning rate and number of epochs. Specifically, we retrain the model (excluding the forget set $\mathcal{D}_{f}$) using the same learning rate and epochs as in the original training for the Retrain method. We limit FT at 40 epochs to mitigate excessive time costs. For PL, GA, RL, and SalUn, we find that strong unlearning can be achieved within 10 epochs on large-scale data. Finally, considering the balance of logit-based and representation-based metrics, we fix the maximum epochs for DUCK, CU, SCAR, and SCRUB at 15, 80, 20, and 90, respectively. These upper limits are determined first under the scenario of Random 100 class unlearning. For Random $k$ class experiments ($k\in\{100,200,300\}$), we choose the actual number of epochs while respecting these maximum values, guided by the H-LR score. For Top $k$ class unlearning, we reuse the hyperparameter configurations established in the corresponding Random $k$ setting. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
@@ Appendix C Additional Experiments
Table 3 provides an extended comparison of the baselines using alternative backbones, specifically ConvNeXt-T [25] and Swin-T [24]. We compare only model-agnostic methods.